Research Report – November 12th

Science Report 12 November 2019
SOL: 2
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: The first 10 samples from the new micrometeorite sampling grid at MM5 (518462E 4248949N) were received, processed and prepared for dispatch. Previously collected gypsum samples were retrieved from storage in preparation for the ISRU cement/concrete manufacture experiments.
Green Hab: Green Hab operations have continued. See Green Hab report.
RAM Operations: N/A
EVA: The second EVA for science saw us returning to MM5 (518462E 4248949N) to complete setting out the micrometeorite sampling grid and collect the first row of 10 samples. See EVA Report.
EVA Suit Maintenance: Suit 10 has had the blown fuse replaced and is ready to be returned to use. Suit 2 is still being evaluated. See Operations Report.
Additional activities: Acoustic volumes were recorded. Habitat layout was evaluated. Psycho-social stressors continued to be monitored.

Copyright © The Mars Society. All rights reserved. | Main Site